# Mean-field approach to finite-size fluctuations in the Kuramoto-Sakaguchi model

### v1

@@ Abstract
The results obtained were verified with simulations for the full KS model and were shown to perform well away from criticality. On the other hand, our approximations were less accurate when approaching the critical coupling strength $K_{\rm{crit}}\approx 1.92$ demarcating the transition from a desynchronized to a partially synchronized state. It seems that near criticality, long-range correlations cannot be neglected and our independence assumptions cease to be valid. We envisage extensions of our framework to the near critical case using a perturbative approach around the expressions derived in this work.
Our approach is sufficiently general to allow for further applications to other interacting particle systems such as $\theta$-neuron models and more complex network topologies going beyond the all-to-all coupling considered in this Letter. The mathematical tools developed here can potentially be adapted to study fluctuations in coupled oscillator systems in the presence of noise [MasKK2010, Luc2014, Gottwald17, PikDG2019, DelgadinoEtAl21, ZagliEtAl23] and finite-size corrections of Lyapunov exponents [CarGP2018].
In summary, our results fully specify the statistical behavior of finite-size fluctuations in the KS model which have been numerically shown to be Gaussian processes. This opens up the way to perform systematic stochastic model reductions for such interacting particle systems from first principles and to construct reduced stochastic equations for designated collective variables [Got2015, SmithGottwald20].
Appendix: Derivations of equations and additional results
Thermodynamic limit for the KS model (1). Most of the dynamical states found in the KS model (1) with finite $N$, have complex chaotic behavior. However, in the thermodynamic limit $N\to\infty$, they admit a simple analytic representation. If $N\gg 1$, the state of the phase oscillators $\{\psi_{i}(t)\}$ can be described by a probability density function $f(\psi,\omega,t)$, which obeys the continuity equation
@@
is the continuum version of the complex order parameter (see formula (3) with $\Omega=0$).
It is well-known [OttA2008, OttA2009] that the long-term dynamics of the continuity equation (28) asymptotically approaches the Ott-Antonsen manifold for $t\to\infty$, consisting of distributions of the form
@@
Remarkably, all statistically stationary states of system (1) lie on the Ott-Antonsen manifold [OmeW2012, OmeW2013]. In particular, the zero solution $z(\omega,t)=0$ of Eq. (30) corresponds to the completely incoherent state in (1), while all stationary partially synchronized states in (1) are represented by rotating waves
$z(\omega,t)=h\left(\frac{\displaystyle\omega-\Omega_{\infty}}{\displaystyle Kr_{\infty}}\right)e^{i\Omega_{\infty}t}$ (31)
@@
The physical meaning of the parameters $r_{\infty}$ and $\Omega_{\infty}$ can be understood by inserting the probability density $f(\psi,\omega,t)$ corresponding to (31) into Eq. (29). This yields
@@
If in the KS model (1) we move into the corotating frame of reference (13), the probability density function in the new frame will take on the time-independent form (14), which is used for the calculations in the main text.
Explicit solutions of the Adler equation. Let us consider the Adler equation written in complex form,
@@
where $z_{*}=i\Delta/H$. (For brevity of notations, we do not specify that $\xi$, $\chi_{1}(t)$, $\chi_{2}(t)$ and $z_{*}$ depend on $\Delta$ and $H$.)
@@
Derivation of Eq. (9) from Eq. (24). Using (5) it is easy to verify that
@@
Now, we consider the integrand of Eq. (24) recalling that $H=KZ_{\infty}e^{-i\lambda}=-iKr_{\infty}e^{i\phi}$. Recognizing that in this case
@@
Pseudo-covariance function for a partially synchronized state. Using the definition of the pseudo-covariance function $\tilde{R}_{\rm{ps}}(t)=\langle\zeta(t)\zeta(t+\tau)\rangle$ and assuming that the variables $e^{i\theta_{j}(t)}$ and $e^{i\theta_{k}(t)}$ are independent for $j\neq k$, by analogy with (23), we obtain
$\tilde{R}_{\mathrm{ps}}(\tau)=\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{i\theta_{j}(t+\tau)}\rangle-\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}\rangle^{2}.$
@@
Details of the numerical setup. Given a distribution $g(\omega)$, we generate a set of $N$ natural frequencies $\omega_{j}$ according to
@@
This equiprobable sampling avoids finite-size effects such as randomly clustered natural frequencies leading to small synchronized clusters with non-zero mean frequencies and their respective interactions [PetP2018, FialkowskiEtAl23]. For each pair of parameters $K$ and $\lambda$, we perform simulations using a standard Runge-Kutta solver with constant time step $dt=0.02$. We start from initial conditions chosen randomly from the interval $0,2\pi]$, discard a transient period of $T_{\mathrm{transient}}=10^{5}$ time units and use the subsequent interval of length $T_{\mathrm{max}}=10^{5}$ time units to calculate statistical quantities such as means, variances, and covariances.
@@
Additional numerical results. To demonstrate how the accuracy of our method changes as we approach the critical coupling strength $K_{\mathrm{crit}}\approx 1.92$, we show additional results for partially synchronized states at $K=5,4,3,2$ for $\lambda=\pi/4$, see Figs. 36, as well as for a completely incoherent state at $K=1$ and $\lambda=\pi/4$, see Fig. 7.

### v2

@@ Abstract
Our theory relies on the assumption of an underlying central limit theorem governing the Gaussian statistics of finite-size fluctuations. Remarkably, the statistical independence of fluctuations of distinct oscillators, which were essential for finding explicit expressions, follows directly from our solution of the Adler equation.
Simulations for the full KS model showed that our theory performs well away from criticality. On the other hand, our approximations were less accurate when approaching the critical coupling strength $K_{\rm{crit}}$, demarcating the transition from a desynchronized to a partially synchronized state. The explicit expressions for the covariance function rely on the explicit solution of the Adler equation. Hence, the quality of our theory crucially depends on the Adler equation being a good approximation of the full system. This is only guaranteed if fluctuations can be neglected in Eq. (1), and breaks down when approaching criticality (see Appendix G for details). We envisage extensions of our framework to the near critical case using a perturbative approach around the expressions derived in this work.
Our approach is sufficiently general to allow for further applications to other interacting particle systems such as $\theta$-neuron models and more complex network topologies going beyond the all-to-all coupling considered in this Letter. The mathematical tools developed here can potentially be adapted to study finite-size fluctuations in noisy coupled oscillator systems [31, 29, 19, 52, 43, 12, 56, 4] finite-size sampling effects [39, 38, 14] and dynamics of complexified KS models [51, 28]. Further, it is known that the incoherent state and partially synchronized states in the KS model are weakly chaotic attractors, where the chaoticity can be attributed to the presence of finite-size fluctuations. Therefore, it is natural to expect that knowing the characteristics of finite-size fluctuations, may aid in computing finite-size corrections of Lyapunov exponents [6].
In summary, our results fully specify the statistical behavior of finite-size fluctuations in the KS model which have been numerically shown to be Gaussian processes. This opens up the way to perform systematic stochastic model reductions for such interacting particle systems from first principles and to construct reduced stochastic equations for designated collective variables [17, 46].
Acknowledgements
The work of O.E.O. was partly supported by the Deutsche Forschungsgemeinschaft under Grant No. OM 99/2-3. O.E.O acknowledges the financial support and hospitality of the Sydney Mathematical Research Institute (SMRI), where parts of this work was completed on an International Visitor Program. The authors are grateful for insightful discussions with A. Pikovsky.
Appendix A Thermodynamic limit for the KS model (1)
Most of the dynamical states found in the KS model (1) with finite $N$, have complex chaotic behavior. However, in the thermodynamic limit $N\to\infty$, they admit a simple analytic representation. If $N\gg 1$, the state of the phase oscillators $\{\psi_{i}(t)\}$ can be described by a probability density function $f(\psi,\omega,t)$, which obeys the continuity equation
@@ Appendix A Thermodynamic limit for the KS model (1)
is the continuum version of the complex order parameter (2).
It is well-known [35, 36] that the long-term dynamics of the continuity equation (27) asymptotically approaches the Ott-Antonsen manifold for $t\to\infty$, consisting of distributions of the form
@@
Remarkably, all statistically stationary states of the KS model (1) lie on the Ott-Antonsen manifold [33, 34]. In particular, the zero solution $z(\omega,t)=0$ of Eq. (29) corresponds to the completely incoherent state in (1), while all stationary partially synchronized states in (1) are represented by rotating waves
$z(\omega,t)=h\left(\frac{\displaystyle\omega-\Omega_{\infty}}{\displaystyle Kr_{\infty}}\right)e^{i\Omega_{\infty}t},$ (30)
@@
The physical meaning of the parameters $r_{\infty}$ and $\Omega_{\infty}$ can be understood by inserting the probability density function $f(\psi,\omega,t)$ corresponding to (30) into Eq. (28). This yields
@@
If in the KS model (1) we move into the corotating frame
$\theta_{i}(t)=\psi_{i}(t)-\Omega_{\infty}t+\phi\quad\mbox{with some}\quad\phi\in\mathbb{R},$ (33)
the probability density function in the new frame will take on the time-independent form (13), which is used for the calculations in the main text. Importantly, for an appropriately chosen value of $\phi$, we can guarantee that (16) yields the thermodynamic limit of the modified order parameter $Z_{\mathrm{mod}}(t)$.
Appendix B Explicit solutions of the Adler equation
Let us consider the Adler equation written in complex form,
@@ Appendix B Explicit solutions of the Adler equation
where $z_{*}=i\Delta/H$. (For brevity of notation, we do not specify that $\xi$, $\chi_{1}(t)$, $\chi_{2}(t)$ and $z_{*}$ depend on $\Delta$ and $H$.)
@@
Appendix C Derivation of Eq. (8) from Eq. (24)
Using (4) it is easy to verify that
@@
Now, we consider the integrand of Eq. (24), recalling that $H=KZ_{\infty}e^{-i\lambda}=-iKr_{\infty}e^{i\phi}$. Recognizing that in this case
@@
[…]
Figure 3: Empirical histograms of the real (a) and imaginary (b) part of the fluctuations $\zeta$ and of $d[{\rm{arg}}Z_{\mathrm{c}}(t)]/dt$ centered around $\Omega_{\infty}\approx-1.71$ (c) for the KS model (1) with $N=1,000$ oscillators, phase lag $\lambda=\pi/4$, and $K=3$. The continuous lines (online red) show a best-fit Gaussian with the same variance.
[…]
Figure 4: Variance of the order parameter $V$ (a) and of the real (b) and imaginary (c) part of the fluctuations $\zeta$ as a function of the non-entrained oscillators $N_{r}$ for the KS model (1) with phase lag $\lambda=\pi/4$ and $K=3$ for $N=60,125,250,500,1000$. The continuous lines (online red) are lines of best fit indicating a scaling law of $N_{r}^{-0.96}$, $N_{r}^{-0.96}$ and $N_{r}^{-0.95}$, respectively.
Appendix D Pseudo-covariance function for a partially synchronized state
Using the definition of the pseudo-covariance function $\tilde{R}_{\rm{ps}}(t)=\langle\zeta(t)\zeta(t+\tau)\rangle$ and employing the established independence of the functions $e^{i\theta_{j}(t)}$ and $e^{i\theta_{k}(t)}$ for $j\neq k$, by analogy with (23), we obtain
$\tilde{R}_{\mathrm{ps}}(\tau)=\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}e^{i\theta_{j}(t+\tau)}\rangle-\frac{1}{N}\sum\limits_{j\in\mathcal{R}}\langle e^{i\theta_{j}(t)}\rangle^{2}.$
@@ Appendix D Pseudo-covariance function for a partially synchronized state
Appendix E Details of the numerical setup
Given a distribution $g(\omega)$, we generate a set of $N$ natural frequencies $\omega_{j}$ according to
@@ Appendix E Details of the numerical setup
This equiprobable sampling avoids finite-size effects such as randomly clustered natural frequencies leading to small synchronized clusters with non-zero mean frequencies and their respective interactions [39, 14]. For each pair of parameters $K$ and $\lambda$, we perform simulations using a standard Runge-Kutta solver with constant time step $dt=0.02$. We start from initial conditions chosen randomly from the interval $0,2\pi]$, discard a transient period of $T_{\mathrm{transient}}=10^{5}$ time units and use the subsequent interval of length $T_{\mathrm{max}}=10^{5}$ time units to calculate statistical quantities such as means, variances, and covariances.
Appendix F Numerical evidence for the assumption on the Gaussianity of fluctuations
Previous work [55, 24] has shown numerically that finite-size fluctuations in the Kuramoto-Sakaguchi model (1) constitute a Gaussian process, consistent with an underlying central limit theorem. For completeness, we provide here further numerical evidence. Fig. 3(a) and (b) show the empirical histogram of the real and the imaginary part of the fluctuations $\zeta$ obtained from a long time simulation over $50,000$ time units with a sampling time of $dt=0.05$ of the KS model together with a best fit Gaussian clearly establishing the Gaussian character of the fluctuations. For $K=3$ we find a value of the skewness for the real and imaginary part of the fluctuations $\zeta$ of 0.058-0.058 and 0.051-0.051, respectively, and a kurtosis of $2.98$ for both. We further show in Fig. 3(c) the empirical histogram for the temporal derivative of the mean phase of the entrained oscillators $\tfrac{d}{dt}{\rm{arg}}Z_{c}(t)$, which also shows a clear Gaussian statistics, consistent with the phase diffusion discussed below.
Figure 5: Normalized covariance $R_{\rm{ps}}(\tau)/R_{\rm{ps}}(0)$ (real part only) of the KS model (1) for partially synchronized states with $N=1,000$ oscillators at $\lambda=\pi/4$ for various values of the coupling strength $K$.
To further establish numerical evidence for an underlying central limit theorem we show in Fig. 4 that the variance of the fluctuations $\zeta/\sqrt{N}$ scales with the number of non-entrained oscillators $N_{r}$ as $1/N_{r}$, as expected for Gaussian fluctuations described by an underlying central limit theorem. We show the scaling for the variance of the order parameter in Fig. 4(a) and of the real and imaginary part of the fluctuations $\zeta$ in Fig. 4 (b) and (c), respectively.
We remark that the central limit theorem holds for all values of the coupling strength $K>K_{\rm{crit}}$. For $K=1.98$ we find a value of the skewness for the real and imaginary part of the fluctuations $\zeta$ of 0.024-0.024 and 0.007-0.007, and a kurtosis of $2.98$ and $3.04$, respectively, suggestive of a Gaussian distribution. The Gaussian character of the empirical probability density function of the fluctuations alone, however, is not sufficient for the existence of a functional central limit theorem which is concerned with convergence to a stochastic process rather than to just random variables. We show in Figure 5 plots of the covariance function of the real part of the fluctuations, $\mathrm{Re}\,R_{\mathrm{ps}}(\tau)$ which decays to zero for all $K>K_{\rm{crit}}$ with the decay of correlation time increasing when approaching criticality.
Appendix G Limitations of the theory
From the previous Section it is clear that in order for the central limit theorem to be valid we require sufficiently large values of $N$. Indeed, for small values of $N=50$ we only have $N_{r}=6$ for $K=4$. Figure 6(a) and (b) clearly show that the empirical histograms of the real and imaginary parts of fluctuations $\zeta$ are far from Gaussian. The speed of the phase of the synchronized order parameter $d[{\rm{arg}}Z_{\mathrm{c}}(t)]/dt$ is still Gaussian, illustrating the phase diffusion discussed below.
[…]
Figure 6: Empirical histograms of the real (a) and imaginary (b) part of the fluctuations $\zeta$ and of $d[{\rm{arg}}Z_{\mathrm{c}}(t)]/dt$ centered around $\Omega_{\infty}\approx-2.39$ (c) for the KS model (1) with $N=50$ oscillators, phase lag $\lambda=\pi/4$, and $K=4$. The continuous lines (online red) show a best-fit Gaussian with the same variance.
Figure 7: Ratio of the standard deviation of the fluctuations $\sqrt{V}$ of the order parameter $r(t)$ and its mean $\langle r(t)\rangle$ for $N=1,000$ oscillators at $\lambda=\pi/4$ as a function of the coupling strength $K$.
The explicit expressions for the covariance functions crucially depends on the explicit solution of the Adler equation (12). To obtain the Adler equation we neglected fluctuations of the modified order parameter, replacing $Z_{\mathrm{mod}}(t)$ by its constant thermodynamic limit. This is only valid provided the fluctuations are small in size compared to the thermodynamic mean values. Approaching criticality, however, the size of fluctuations increases. Figure 7 shows the ratio of the standard deviation of the fluctuations $\sqrt{V}$ of the order parameter $r(t)$ and its mean as a function of the coupling strength $K$. It is clearly seen how fluctuations increase rapidly in size when approaching criticality at $K_{\rm{crit}}\approx 1.92$. This implies that the Adler equation and its solution become less valid closer to criticality. Hence, the quality of our theory depends on the value of the coupling strength $K$. Indeed, Figure 2(c) in the main text illustrates that our theory works well for coupling strengths sufficiently far from criticality, and gets worse the closer we approach criticality.
Appendix H Additional numerical results
@@
To demonstrate how the accuracy of our method changes as we approach the critical coupling strength $K_{\mathrm{crit}}\approx 1.92$, we show in Figs. 811 additional results for partially synchronized states at $K=5,4,3,2$ for $\lambda=\pi/4$, and in Fig. 12 additional results for a completely incoherent state at $K=1$ and $\lambda=\pi/4$. We plot results from numerical simulations of the the KS model (1) and of our theory, Eq. (8) for the partially synchronized case and Eq. (10) for the incoherent state, respectively.
@@
In addition, to demonstrate the effectiveness of our theoretical formulas for other system parameters, we present results for different values of the phase lag $\lambda$ and different frequency distributions $g(\omega)$. Fig. 13 shows how the variance $V$ of order parameter depends on the coupling strength $K$ for a unit variance Gaussian distribution (top row), for a triangular distribution (middle row)
$g(\omega)=\left\{\begin{array}[]{lcl}1-|\omega|&\mbox{ if }&|\omega|\leq 1,\\[2.84526pt] 0&\mbox{ if }&|\omega|>1,\end{array}\right$.
and for a Lorentzian distribution (bottom row)
$g(\omega)=\frac{1}{\pi(1+\omega^{2})},$
for several values of the phase lag $\lambda$. It is seen that the quality of our approximations does not significantly depend on neither the frequency distribution nor on the value of the phase lag.
Figure 13: Variance $V$ of the order parameter $r(t)$ as a function of the coupling strength $K$ for different values of the phase lag $\lambda$ for three frequency distributions $g(\omega)$: Gaussian (a)–(c), Triangular (d)–(f), and Lorentzian (g)–(i). Numerical simulations for the KS model (1) with $N=1,000$ (blue squares) vs. theoretical predictions (9) and (11) (solid curves).
Finally, in Figure 14 we show that the results for the full KS model (1) are indistinguishable by eye when a larger network with $10,000$ oscillators is simulated, indicative that $N=1,000$ are sufficient to study the thermodynamic limit.
Figure 14: Comparison of the results from numerical simulations of the KS model (1) for $N=1,000$ and $N=10,000$. (a)-(b): The solid curves show the covariance $R_{\mathrm{ps}}(\tau)$ of fluctuations $\zeta(t)$ of the modified order parameter $Z_{\mathrm{mod}}(t)$ for a partially synchronized state at $K=6$ and $\lambda=\pi/4$. Dashed curves show the theoretical prediction (8). (c): Variance of the order parameter $V$ as a function of the coupling strength $K$ (squares and crosses). The theoretical predictions (8) and (9) are shown as solid curves.
Appendix I Phase diffusion
Mean-field theory for partially synchronized states predicts that the argument of the complex order parameter
$Z(t)=\frac{1}{N}\sum\limits_{j=1}^{N}e^{i\psi_{j}(t)},$
uniformly increases or decreases with time. To subtract this drift, we can move into the corotating frame of reference and consider $\arg(Z(t)e^{-i\Omega t})$, where $\Omega$ is the average speed. However, this quantity experiences non-stationary irregular behaviour, resembling Brownian motion, as shown in Fig. 15. This phenomenon is known as phase diffusion [42, 40]. Due to this phase diffusion, the function $Z(t)e^{-i\Omega t}$ is difficult to characterize. Therefore, in our study we work with a modified order parameter
$Z_{\mathrm{mod}}(t)=Z(t)\frac{\overline{Z_{\mathrm{c}}(t)}}{|Z_{\mathrm{c}}(t)|},$
where
$Z_{\mathrm{c}}(t)=\frac{1}{N}\sum\limits_{j\in\mathcal{C}}e^{i\psi_{j}(t)}$
is the order parameter, defined as the sum over the coherent (locked) oscillators only. In numerical simulations, we identify the $j$th oscillator as coherent if its natural frequency $\omega_{j}$ satisfies the locking condition $|\omega_{j}-\Omega|\leq K\langle|Z(t)|\rangle$. It turns out, that the argument of the modified order parameter $Z_{\mathrm{mod}}(t)$ does not exhibit phase diffusion phenomenon, and the behavior of this complex function is nicely approximated by a two-dimensional Gaussian process [55]. This numerical observation is crucial for our analytical approach.
Figure 15: Phase diffusion in the KS model (1). The argument of the complex order parameter in the uniformly corotating frame (black curve) exhibits non-stationary random walk behavior. Contrary, the argument of the modified order parameter $Z_{\mathrm{mod}}(t)$ (red curve) exhibits stationary statistical behaviour. Shown are numerical results for the KS model (1) with $N=1,000$, $K=2$, $\lambda=\pi/4$, and a unit width Gaussian frequency distribution $g(\omega)$.
Appendix J Pairwise covariances of rogue oscillators
In Proposition 1, we derived an explicit formula for the solutions of the Adler equation (34). Simple calculations show that this formula has the equivalent form
$e^{i\theta(t)}e^{-i\theta(0)}=\frac{\displaystyle\chi_{+}e^{i\xi t}-\chi_{-}}{\displaystyle-\overline{\chi}_{-}e^{i\xi t}+\overline{\chi}_{+}}=\frac{\displaystyle\chi_{+}-\chi_{-}e^{-i\xi t}}{\displaystyle-\overline{\chi}_{-}+\overline{\chi}_{+}e^{-i\xi t}},$
where
$\chi_{-}=iHe^{-i\theta(0)}-(\Delta-\xi),\quad\chi_{+}=iHe^{-i\theta(0)}-(\Delta+\xi).$
We use this form to justify the following statements.
Proposition 2
If $\Delta>|H|$, we have $|\chi_{-}/\chi_{+}|<1$ and the following expansion holds
$e^{i\theta(t)}$ $=$ $-\frac{\displaystyle\chi_{-}}{\displaystyle\overline{\chi}_{+}}e^{i\theta(0)}+\frac{\displaystyle|\chi_{+}|^{2}-|\chi_{-}|^{2}}{\displaystyle\overline{\chi}_{+}\overline{\chi}_{-}}e^{i\theta(0)}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle\overline{\chi}_{-}}{\displaystyle\overline{\chi}_{+}}\right)^{n}e^{in\xi t}$ (38)
$=$ $\left\langle e^{i\theta(t)}\right\rangle-\frac{\displaystyle 2i\xi}{\displaystyle\overline{H}}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta-\xi}{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta+\xi}\right)^{n}e^{in\xi t},$
where $\left\langle e^{i\theta(t)}\right\rangle$ denotes the time average of $e^{i\theta(t)}$. On the other hand, if $\Delta<-|H|$, then $|\chi_{+}/\chi_{-}|<1$ and
$e^{i\theta(t)}$ $=$ $-\frac{\displaystyle\chi_{+}}{\displaystyle\overline{\chi}_{-}}e^{i\theta(0)}+\frac{\displaystyle|\chi_{-}|^{2}-|\chi_{+}|^{2}}{\displaystyle\overline{\chi}_{+}\overline{\chi}_{-}}e^{i\theta(0)}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle\overline{\chi}_{+}}{\displaystyle\overline{\chi}_{-}}\right)^{n}e^{-in\xi t}$
$=$ $\left\langle e^{i\theta(t)}\right\rangle+\frac{\displaystyle 2i\xi}{\displaystyle\overline{H}}\sum\limits_{n=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta+\xi}{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta-\xi}\right)^{n}e^{-in\xi t}.$
Proof: For brevity, we denote $\zeta=iHe^{-i\theta(0)}$. Then $e^{i\theta(0)}=i\overline{\zeta}/\overline{H}$. Next, using the definition $\xi=\sqrt{\Delta^{2}-|H|^{2}}$, we obtain
$|\chi_{+}|^{2}$ $=$ $(\zeta-\Delta-\xi)(\overline{\zeta}-\Delta-\xi)$
$=$ $(\Delta+\xi)(2\Delta-\zeta-\overline{\zeta}),$
$|\chi_{-}|^{2}$ $=$ $(\zeta-\Delta+\xi)(\overline{\zeta}-\Delta+\xi)$
$=$ $(\Delta-\xi)(2\Delta-\zeta-\overline{\zeta}),$
and
$\left(|\chi_{+}|^{2}-|\chi_{-}|^{2}\right)e^{i\theta(0)}$ $=$ $2\xi(2\Delta-\zeta-\overline{\zeta})i\overline{\zeta}/\overline{H}$
$=$ $-2i\xi(\overline{\zeta}^{2}-2\Delta\overline{\zeta}+|H|^{2})/\overline{H},$
$\overline{\chi}_{+}\overline{\chi}_{-}$ $=$ $(\overline{\zeta}-\Delta-\xi)(\overline{\zeta}-\Delta+\xi)$
$=$ $\overline{\zeta}^{2}-2\Delta\overline{\zeta}+|H|^{2}.$
Therefore,
$\frac{\displaystyle|\chi_{+}|^{2}-|\chi_{-}|^{2}}{\displaystyle\overline{\chi}_{+}\overline{\chi}_{-}}e^{i\theta(0)}=-\frac{\displaystyle 2i\xi}{\displaystyle\overline{H}}.$
On the other hand, we have
$\frac{\displaystyle\overline{\chi}_{-}}{\displaystyle\overline{\chi}_{+}}=\frac{\displaystyle\overline{\zeta}-\Delta+\xi}{\displaystyle\overline{\zeta}-\Delta-\xi}=\frac{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta-\xi}{\displaystyle i\overline{H}e^{i\theta(0)}+\Delta+\xi}.$
Summarizing, this gives formula (38). The other formula in the proposition can be checked similarly.
Proposition 3
Suppose that $\theta_{j}(t)$ and $\theta_{k}(t)$ are two solutions of Eq. (34) with $\Delta=\Delta_{j}$ and $\Delta=\Delta_{k}$, respectively. The complex parameter $H$ in Eq. (34) is assumed to be the same for both indices $j$ and $k$. If $|\Delta_{j}|>|H|$ and $|\Delta_{k}|>|H|$, then
$\hskip-19.91692pt\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle-\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{-i\theta_{k}(t)}\rangle=$
$=\frac{\displaystyle 4\xi_{j}\>\mathrm{sgn}\>\Delta_{j}\>\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle|H|^{2}}\times$
$\times\sum\limits_{n,m=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}-\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}+\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}\right)^{n}\times$
$\times\left(\frac{\displaystyle iHe^{-i\theta_{k}(0)}-\Delta_{k}+\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle iHe^{-i\theta_{k}(0)}-\Delta_{k}-\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}\right)^{m}\times$
$\times e^{-im\xi_{k}\tau\>\mathrm{sgn}\>\Delta_{k}}\delta(n\xi_{j}\>\mathrm{sgn}\>\Delta_{j}-m\xi_{k}\>\mathrm{sgn}\>\Delta_{k}),$
and
$\hskip-19.91692pt\langle e^{i\theta_{j}(t)}e^{i\theta_{k}(t+\tau)}\rangle-\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{i\theta_{k}(t)}\rangle=$
$=-\frac{\displaystyle 4\xi_{j}\>\mathrm{sgn}\>\Delta_{j}\>\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle\overline{H}^{2}}\times$
$\times\sum\limits_{n,m=1}^{\infty}\left(\frac{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}-\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}{\displaystyle i\overline{H}e^{i\theta_{j}(0)}+\Delta_{j}+\xi_{j}\>\mathrm{sgn}\>\Delta_{j}}\right)^{n}\times$
$\times\left(\frac{\displaystyle i\overline{H}e^{i\theta_{k}(0)}+\Delta_{k}-\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}{\displaystyle i\overline{H}e^{i\theta_{k}(0)}+\Delta_{k}+\xi_{k}\>\mathrm{sgn}\>\Delta_{k}}\right)^{m}\times$
$\times e^{im\xi_{k}\tau\>\mathrm{sgn}\>\Delta_{k}}\delta(n\xi_{j}\>\mathrm{sgn}\>\Delta_{j}+m\xi_{k}\>\mathrm{sgn}\>\Delta_{k}),$
where $\xi_{j}=\sqrt{\Delta_{j}^{2}-|H|^{2}}$ and $\xi_{k}=\sqrt{\Delta_{k}^{2}-|H|^{2}}$.
Proof: The above formulas are straightforward consequences of Proposition 2 and the identity $\langle e^{i\nu t}\rangle=\delta(\nu)$
Remark 4
If $\xi_{j}$ and $\xi_{k}$ in Proposition 3 are incommensurable, then
$\langle e^{i\theta_{j}(t)}e^{-i\theta_{k}(t+\tau)}\rangle=\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{-i\theta_{k}(t+\tau)}\rangle$
and
$\langle e^{i\theta_{j}(t)}e^{i\theta_{k}(t+\tau)}\rangle=\langle e^{i\theta_{j}(t)}\rangle\>\langle e^{i\theta_{k}(t+\tau)}\rangle.$
Appendix K Daido’s theory